26. In the process of producing 1 hormone-free, stress-free, happy cow, the New Age Cattle Ranch produces I ton of natural, organic plant fertilizer. Demand for the ranchs cattle is given by Pe -2,000 5Q, and demand for fertilizer is given by P 500 4Q. The cost of producing Q cow-fertilizer bundles is C- 5,000+4000. To maximize profits, the ranch should a. produce 160 cow-fertilizer units, sell 160 cattle, and sell 62.50 tons of fertilizer. b. produce 160 cow-fertilizer units, sell 160 cattle, and sell 12.50 tons of fertilizer. c. produce 160 cow-fertilizer units, sell 160 cattle, and sell 160 tons of fertilizer. d. produce 116.67 cow-fertilizer units, sell 116.67 cattle, and sell 116.67 tons of fertilizer. e. produce 200 cow-fertilizer units, sell 200 cattle, and dispose of all the fertilizer in an environmentally friendly manner.
Answer #1

Given, P, 2000-50.

Thus, total revenue will be

large TR_c=P_c Q_c= (2000 - 5Q_c)Q_c

TR = 2000Q.. _ 5

Now differentiate wrt Qc we get

R2000-100

Similarly, the fertilizer Marginal revenue can be determined

P 500-4Q

large TR_f = 500Q_f - 4Q_f^2

Differentiating above equation wrt Qf we get

large MR_f = 500 - 8Q_f

The total cost function of the firm is given by

C = 5000 + 400Q

MC = 400

In case of cow equate MR WITH MC

2000100400

10G.-2000-400

160

P, 2000-5 * 160-2000-800-$1,200

Similarly equate fertilizers MR to MC we get

500- 8f 400

500 - 400

100 12.5unit.s =-=

Pf 500 - 4Q 500 412.5 450perunit

Option B is correct.

Produce 160 cow-fertilizer units, sell 160 and sell 12.5 tons of fertilizer.

Please contact if having any query thank you.